BIBTEX(1) BIBTEX(1) NAME bibtex−makea(La)TeXbibliography SYNOPSIS bibtexauxname DESCRIPTION BibTeXreadsthetop-levelauxiliary(.aux)filethatwasoutputduringtherunningof(la)tex(1) andcreatesabibliography(.bbl)filethatwillbeincorporatedintothedocumentonsubsequent runsof(La)TeX.Theauxnameonthecommandlinemustbegivenwithoutthe.auxextension. BibTeXlooksup,inthebibliographicdatabase(.bib)files,theentriesspecifiedbythe\citeand \nocitecommandsinthe(La)TeXsourcefile.Itformatstheinformationfromthoseentries accordingtoinstructionsinthestyle(.bst)file,anditoutputstheresultstothe.bblfile. TheLaTeXmanual(akaLaTeX-ADocumentPreparationSystem,byLeslieLamport)explains whataLaTeXsourcefilemustcontaintoworkwithBibTeX.AppendixBoftheLaTeXmanual describestheformatofthe.bibfiles.The"BibTeXing"documentdescribesextensionsanddetails ofthatformat,anditgivesotherusefulhintsforusingBibTeX. ENVIRONMENT bibtexsearchesthedirectoriesinthepathdefinedbytheTEXINPUTSenvironmentvariablefor .bstfiles.IfTEXINPUTSisnotset,itusesthesystemdefault:.:/usr/TeX/lib/tex/macros//. For.bibfiles,itusestheBIBINPUTSenvironmentvariableinasimilarway.Thedefaultpath thereis:.:/usr/TeX/lib/tex/bib.Seetex(1)forthedetailsofthesearching.IftheenvironmentvariableTEXMFOUTPUTisset,BibTeXattemptstoputitsoutputfilesinit,iftheycannot beputinthecurrentdirectory.Again,seetex(1).Nospecialsearchingisdoneforthe.auxfile. FILES btxdoc.tex"BibTeXing"−LaTeXabledocumentationforgeneralBibTeXusers btxhak.tex"DesigningBibTeXStyles"−LaTeXabledocumentationforstyledesigners btxmac.texmacrosforusingBibTeXwithplainTeX btxdoc.bibdatabasefileforthosetwodocuments xampl.bibdatabasefilegivingexamplesofallstandardentrytypes btxbst.doctemplatefileanddocumentationforthestandardstyles Allthosefilesshouldbeavailablesomewhereonthesystem. SEEALSO latex(1),tex(1). 1